# Preserving curvature lower bounds when Ricci flowing non-smooth initial data

### v1

@@ 3. $C^{0}$ Limits of smooth Riemannian manifolds with curvature lower bounds
Let $t_{0}$ be a time with $t_{0}\leq\hat{T}(n,h,\hat{\varepsilon})$ small, $\hat{T}(n,h,\hat{\varepsilon})$ to be specified below.
We scale the solution, by $\hat{g}(\hat{t})=cg(\frac{\hat{t}}{c})$ and $\hat{h}=ch$ with $c=\frac{1}{t_{0}}$. If $t_{0}\leq\hat{T}(n,h,\hat{\varepsilon})$ is chosen small enough, then we have $\sum_{j=0}^{4}|{{}^{\hat{h}}\nabla}^{j}{\rm Rm}(\hat{h})|^{2}\leq\hat{\varepsilon}.$ Assume for the moment that $inj_{h}(x)\geq 1$ for all $x\in B_{\hat{h}}(p,1).$ Then there exist geodesic coordinates on a ball of radius one centred at $p$ such that $|\hat{h}-\delta|+|D\hat{h}|+|D^{2}\hat{h}|\leq\hat{\varepsilon}$ in these coordinates. The proof of the estimates in [Sim01] and [Shi89] is still valid for $\hat{g}$ and we obtain $\hat{t}|{{}^{\hat{h}}\nabla}\hat{g}|^{2}(\cdot,\hat{t})+\hat{t}^{2}|{{}^{\hat{h}}\nabla}^{2}\hat{g}|^{2}(\cdot,\hat{t})|\leq 1$ for all $\hat{t}\leq 1$. Using these estimates for $\hat{t}=1$, with the estimates for $\hat{h}$ in geodesic coordinates and the fact that $\frac{1}{(1+2\hat{\varepsilon})}\hat{h}\leq\hat{g}(t)\leq(1+2\hat{\varepsilon})\hat{h}$ for all $\hat{t}\leq 1$ (a scale invariant fact), we see, by interpolating, that $|{\rm Rm}(\hat{g})(p,1)|\leq\psi(\hat{\varepsilon})$. This implies $|{\rm Rm}(g)(p,t_{0})|\leq\frac{\psi(\hat{\varepsilon})}{t_{0}}$ as required.
@@
Proof of Theorem 3.7 and Remark 3.8 By using Shi’s Theorem (Theorem 1.2 of [Shi89]) applied to one of the metrics $\hat{h}:=g_{k}(0)$ we can find a metric $h$ with $\sup_{M}|\nabla^{j}{\rm Rm}(h)|^{2}<\infty,$ for all $j\in\mathbb{N}_{0},$ $inj(M,h)\geq i_{0}>0$ and $|g_{i}(0)-h|\leq\hat{\varepsilon},$ for all $i\in\mathbb{N}$, after taking a subsequence if necessary, where $\hat{\varepsilon}$ as small as we like. Hence without loss of generality, $|g_{i}(0)-h|_{h}\leq\hat{\varepsilon}\leq\varepsilon(n)$ for all $i\in\mathbb{N}$, where $\hat{\varepsilon}>0$ is a constant as small as we like. Let $g_{i}(t)_{t\in[0,T(n,h)]}$ be the Ricci DeTurck $h$-flow solution from Theorem 3.1 with initial value $g_{i}(0)$. By choosing $\hat{\varepsilon}$ small enough, and reducing $T$ if necessary, we have, $\sup_{M}|{\rm Rm}(g_{i}(t))|\leq\frac{\sigma(n)}{t}$ for all $t\in(0,T)$ for all $i\in\mathbb{N},$ due to Theorem 3.5, where $\sigma(n)$ is a small constant of our choosing. We choose $S=T/2$ and define $\Phi_{i}(x,S)=x,$ and let $\Phi_{i}:M\times[0,T]\to M$ be the solution to (3.1). Then $\ell_{i}(t)=(\Phi_{i})(\cdot,t)^{*}(g_{i}(t))_{t\in(0,T]}$ is a smooth solution to Ricci flow, in view of Theorem 3.4. Furthermore, Theorem 3.4, tells us that $\Phi_{i}(t)\to\Phi_{i}(0)$ locally uniformly as $t\searrow 0,$ where $\Phi_{i}(0):M\to M$ is a homeomorphism. In this instance the convergence is locally smooth and $\Phi_{i}(0)$ is a diffeomorphism, in view of the fact that $g_{i}(0)$ is smooth. Hence $\ell_{i}(t)=(\Phi_{i}(t))^{*}g_{i}(t)$ is a smooth solution to Ricci flow with initial smooth value $\ell_{i}(0):=(\Phi_{i})^{*}(g_{i}(0))$, and satisfies $|{\rm Rm}(\ell_{i}(t))|\leq\frac{\sigma}{t}$ and $\sup_{M}|{\rm Rm}(\ell_{i}(0))|<\infty$. Theorem 1.2 of Lee’s paper, [Lee19] (see also Simon [Sim08], Chen [BLChen09]), now guarantees that $(M^{n},\ell_{i}(t))_{t\in[0,T]}$ is a bounded curvature solution.
@@ 4. Non-collapsed limits of smooth spaces with curvature lower bounds
Then for $0\leq\varepsilon\leq\frac{1}{100}$ the following holds:
@@
    ${\rm Ric}(g(0))\geq-\varepsilon{\rm R}(g(0))g(0),$
    $\implies$ ${\rm Ric}(g(t))\geq-\varepsilon(1+4t)g(t)-\varepsilon(1+4t){\rm R}(g(t))g(t)$ for all $t\in[0,T]\cap[0,\frac{1}{8}]$
@@
    $\sec(g(0))\geq-\varepsilon$
    $\implies$ $\sec(g(t))\geq-\varepsilon(1+4t)-\varepsilon(1+4t){\rm R}(g(t))$ for all $t\in[0,T]\cap[0,\frac{1}{8}]$
@@
    ${\rm Ric}(g(0))\geq-\varepsilon g(0)$
    $\implies$ ${\rm Ric}(g(t))\geq-\varepsilon(1+100t)g(t)-\varepsilon(1+100t)t{\rm R}(g(t))g(t)$ for all $t\in[0,T]\cap[0,\frac{1}{200}]$
@@
    $\sec(g(0))\geq-\varepsilon$
    $\implies$ $\sec(g(t))\geq-\varepsilon(1+100t)-\varepsilon(1+100t)t{\rm R}(g(t))$ for all $t\in[0,T]\cap[0,\frac{1}{200}]$
@@
The proof of the Lemma involves examining the evolution of Tensors/ geometric inequalities under the Ricci flow and using the Maximum principle of [Ham82, Ham86]. For example in case (c), on a closed manifold, one first writes the evolution equation of the tensor $T(\cdot,t)={\rm Ric}(g(t))+\varepsilon(1+100t)g(t)+\varepsilon(1+100t)t{\rm R}(g(t))g(t)$ as $\frac{\partial}{\partial t}T=\Delta_{g}T+L$ : with the help of the evolution equations for the Ricci curvature and scalar curvature computed in [Ham82], one can explicitly calculate $L$. Then, using the Maximum Principle of Hamilton, one shows that $T(\cdot,0)>0\implies T(\cdot,t)>0.$ The term in $L$ coming from $(\frac{\partial}{\partial t}-\Delta_{g(t)})(t{\rm R}(g(t))),$ is a non-negative term, and plays an important role in the maximum principle argument : It guarantees that $L$ will be positive at any potential first time $t_{0}>0$ and point $p_{0}$ and non-zero direction $v_{0}$ where $T(p_{0},t_{0})(v_{0},v_{0})=0$$\Box$.□ .
@@
Assume for some $\varepsilon\in[0,\frac{1}{100}]$ that ${\rm Ric}(g(0))\geq-\varepsilon g(0)$ $(\sec(g(0))\geq-\varepsilon)$ and that the solution satisfies $|{\rm R}(g(t))|\leq\frac{c^{2}_{0}}{t}.$ Then ${\rm Ric}(g(t))\geq-\varepsilon(1+c^{2}_{0})(1+100t)g(t)$ $(\sec(g(t))\geq-\varepsilon(1+c^{2}_{0})(1+100t))$
@@
iv) In the case that we have merely a negative-bound from below at time zero, for example $\mathcal{R}(g(0))+kI\geq 0$ for a $k>0$, attempts to generalise the above "Ingredient 2" to higher dimensions by adding multiples of $t{\rm R}(g(t))I$ and $I$ to $\mathcal{R}(g(t))$ for example, have, to this date, not been successful, except in the case that one already knows $|{\rm Rm}(g(t))|\leq\frac{c^{2}_{0}}{t}$ and $c^{2}_{0}\leq\varepsilon(n),$ with $\varepsilon(n)$ sufficiently small, as explained in Section 3. A break through for the case that the bound from below is negative in higher dimensions was achieved by Bamler/Cabezas/Rivas [BCRW19]. There they examine the quantity $\ell(t)=\inf\{\alpha\in[0,\infty)\ |\ \mathcal{R}(g(t))+\alpha I\in\mathcal{C}\}$ where $\mathcal{C}$ represents the curvature condition (cone) being considered. Due to some scaling arguments (as in the proof of Theorem 4.5 above), it suffices to consider the case that one has at time zero $\ell(0)\leq\varepsilon(n,v_{0}),$ where $\varepsilon(n,v_{0})>0$ is as small as we like. They show, that the reaction equation of $\ell(t)$ satisfies an inequality of the type
@@
in the setting that $|{\rm Rm}(g(t))|\leq\frac{c^{2}_{0}}{t}$ which then allows them to obtain a bound on $\ell(t)$ by comparing $\ell$ to the solution $u$ to (4.3) whose initial value is constant, $u(0):=\varepsilon$.
@@
In some cases the initial manifold being considered may have some bound on curvature from below, but no bound on the norm of the curvature. Hence even starting the Ricci flow presents difficulties. We survey here some results on the existence of a Ricci flow in this setting. As a prototype case, we consider a three dimensional smooth, complete Riemannian manifold $(M^{3},g(0))$ which is uniformly non-collapsed, that is ${\rm vol}(B_{g(t)}(x,1))\geq v_{0}>0$ for all $x\in M,$ ${\rm Ric}(g(0))\geq-\varepsilon,$ $\varepsilon\leq\varepsilon(v_{0}),$ but we do not assume bounded curvature. That is $\sup_{M}|{\rm Rm}(g(0))|=\infty$ is possible.
@@
In the process of producing the local Ricci flow, it was noticed that the arguments are purely local. These methods were subsequently pushed even further to obtain so called Pyramid Ricci flows, $g(t)$ which are defined on space time regions $\Omega\subseteq M\times(0,T),$ $\Omega:=\cup_{i=1}^{\infty}B_{g(0)}(p,i)\times[0,T_{i}]$ where $T_{1}(v_{0},k)>T_{2}(v_{0},k)>T_{3}(v_{0},k)\ldots$ and $T_{i}\to 0$ as $i\to\infty,$ under the assumption that ${\rm vol}_{g(0)}(B_{g(0)}(p,1))\geq v_{0},$ and $\mathcal{R}(g(0))+kI\in\mathcal{C}$ where $\mathcal{C}$ is one of the conditions (i)-(iv) or (vi)-(ix), and $p\in M$ is a fixed point. As an application, one can prove for example, that any possible Gromov-Hausdorff pointed limit of smooth three manifolds $(M^{3}_{i},g_{i}(0),p_{i})$ satisfying ${\rm vol}_{g_{i}(0)}(B_{g_{i}(0)}(p_{i},1))\geq v_{0},$ and ${\rm Ric}(g_{i}(0))+kg_{i}(0)\geq 0$ is a topological manifold : Simon/Topping [ST21]. See the following papers for further results of this type and further information on this setting: [ST21], McLeod/Topping [MT22], McLeod/Topping [MT20], McLeod [Mcl18], [Hoc19], [LT21b], (cf. Liu [Liu18],[Liu15]).
@@
(Chan/Chen/Lee, Theorem 1.4 in [CCL22]) For any positive integer $n\geq 3$ and constant $A\geq 1000n$, there exists a constant $\varepsilon_{0}(n,A)>0$ such that the following holds. Suppose $(M_{i}^{n},g_{i}(0),p_{i})$ is a pointed sequence of smooth, complete, connected Riemannian manifolds with bounded curvature satisfying the following properties:
@@
where $\nu$ is the localised entropy, a localisation of Perleman’s entropy introduced in [Wan18]. Then there exists solutions $(M_{i},g_{i}(t)),p_{i})_{t\in[0,T)}$ for a uniform $T>0$ (independent of $i$) satisfying, $|{\rm Rm}(g_{i}(t)|\leq\frac{C(n,A)\varepsilon_{0}}{t},$ $inj(M_{i},g_{i}(t))\geq\gamma(n)\sqrt{t}$ for all $t\in(0,T)$ and, after taking a subsequence, the pointed Cheeger-Hamilton limit $(M,g(t),p)_{t\in(0,T)}$ of $(M_{i},g_{i}(t)),p_{i})_{t\in(0,T)}$ exists and also satisfies $|{\rm Rm}(g(t)|\leq\frac{C(n,A)\varepsilon_{0}}{t},$ $inj(M,g(t))\geq\gamma(n)\sqrt{t}$ and
$(\int_{B_{g(t)}(q,2)}|{\rm Rm}(g(t)|^{\frac{n}{2}}d{\rm vol}_{g(t)})^{\frac{2}{n}}\leq C(n,A)\varepsilon_{0},$

### v2

@@ 3. $C^{0}$ Limits of smooth Riemannian manifolds with curvature lower bounds
Let $t_{0}$ be a time with $t_{0}\leq\hat{T}(n,h,\hat{\epsilon})$ small, $\hat{T}(n,h,\hat{\epsilon})$ to be specified below.
We scale the solution, by $\hat{g}(\hat{t})=cg(\frac{\hat{t}}{c})$ and $\hat{h}=ch$ with $c=\frac{1}{t_{0}}$. If $t_{0}\leq\hat{T}(n,h,\hat{\epsilon})$ is chosen small enough, then we have $\sum_{j=0}^{4}|{{}^{\hat{h}}\nabla}^{j}{\rm Rm}(\hat{h})|^{2}\leq\hat{\epsilon}.$ Assume for the moment that $inj_{h}(x)\geq 1$ for all $x\in B_{\hat{h}}(p,1).$ Then there exist geodesic coordinates on a ball of radius one centred at $p$ such that $|\hat{h}-\delta|+|D\hat{h}|+|D^{2}\hat{h}|\leq\hat{\epsilon}$ in these coordinates. The proof of the estimates in [Sim01] and [Shi89] is still valid for $\hat{g}$ and we obtain $\hat{t}|{{}^{\hat{h}}\nabla}\hat{g}|^{2}(\cdot,\hat{t})+\hat{t}^{2}|{{}^{\hat{h}}\nabla}^{2}\hat{g}|^{2}(\cdot,\hat{t})|\leq 1$ for all $\hat{t}\leq 1$. Using these estimates for $\hat{t}=1$, with the estimates for $\hat{h}$ in geodesic coordinates and the fact that $\frac{1}{(1+2\hat{\epsilon})}\hat{h}\leq\hat{g}(t)\leq(1+2\hat{\epsilon})\hat{h}$ for all $\hat{t}\leq 1$ (a scale invariant fact), we see, by interpolating, that $|{\rm Rm}(\hat{g})(p,1)|\leq\psi(\hat{\epsilon})$. This implies $|{\rm Rm}(g)(p,t_{0})|\leq\frac{\psi(\hat{\epsilon})}{t_{0}}$ as required.
@@
Proof of Theorem 3.7 and Remark 3.8 By using Shi’s Theorem (Theorem 1.2 of [Shi89]) applied to one of the metrics $\hat{h}:=g_{k}(0)$ we can find a metric $h$ with $\sup_{M}|\nabla^{j}{\rm Rm}(h)|^{2}<\infty,$ for all $j\in\mathbb{N}_{0},$ $inj(M,h)\geq i_{0}>0$ and $|g_{i}(0)-h|\leq\hat{\epsilon},$ for all $i\in\mathbb{N}$, after taking a subsequence if necessary, where $\hat{\epsilon}$ as small as we like. Hence without loss of generality, $|g_{i}(0)-h|_{h}\leq\hat{\epsilon}\leq\epsilon(n)$ for all $i\in\mathbb{N}$, where $\hat{\epsilon}>0$ is a constant as small as we like. Let $g_{i}(t)_{t\in[0,T(n,h)]}$ be the Ricci DeTurck $h$-flow solution from Theorem 3.1 with initial value $g_{i}(0)$. By choosing $\hat{\epsilon}$ small enough, and reducing $T$ if necessary, we have, $\sup_{M}|{\rm Rm}(g_{i}(t))|\leq\frac{\sigma(n)}{t}$ for all $t\in(0,T)$ for all $i\in\mathbb{N},$ due to Theorem 3.5, where $\sigma(n)$ is a small constant of our choosing. We choose $S=T/2$ and define $\Phi_{i}(x,S)=x,$ and let $\Phi_{i}:M\times[0,T]\to M$ be the solution to (3.1). Then $\ell_{i}(t)=(\Phi_{i})(\cdot,t)^{*}(g_{i}(t))_{t\in(0,T]}$ is a smooth solution to Ricci flow, in view of Theorem 3.4. Furthermore, Theorem 3.4, tells us that $\Phi_{i}(t)\to\Phi_{i}(0)$ locally uniformly as $t\searrow 0,$ where $\Phi_{i}(0):M\to M$ is a homeomorphism. In this instance the convergence is locally smooth and $\Phi_{i}(0)$ is a diffeomorphism, in view of the fact that $g_{i}(0)$ is smooth. Hence $\ell_{i}(t)=(\Phi_{i}(t))^{*}g_{i}(t)$ is a smooth solution to Ricci flow with initial smooth value $\ell_{i}(0):=(\Phi_{i})^{*}(g_{i}(0))$, and satisfies $|{\rm Rm}(\ell_{i}(t))|\leq\frac{\sigma}{t}$ and $\sup_{M}|{\rm Rm}(\ell_{i}(0))|<\infty$. Theorem 1.2 of Lee’s paper, [Lee19] (see also Simon [Sim08], Chen [BLChen09]), now guarantees that $(M^{n},\ell_{i}(t))_{t\in[0,T]}$ is a bounded curvature solution.
@@ 4. Non-collapsed limits of smooth spaces with curvature lower bounds
Then for $0\leq\epsilon\leq\frac{1}{100}$ the following holds:
@@
    ${\rm Ric}(g(0))\geq-\epsilon{\rm R}(g(0))g(0),$
    $\implies$ ${\rm Ric}(g(t))\geq-\epsilon(1+4t)g(t)-\epsilon(1+4t){\rm R}(g(t))g(t)$ for all $t\in[0,T]\cap[0,\frac{1}{8}]$
@@
    $\sec(g(0))\geq-\epsilon$
    $\implies$ $\sec(g(t))\geq-\epsilon(1+4t)-\epsilon(1+4t){\rm R}(g(t))$ for all $t\in[0,T]\cap[0,\frac{1}{8}]$
@@
    ${\rm Ric}(g(0))\geq-\epsilon g(0)$
    $\implies$ ${\rm Ric}(g(t))\geq-\epsilon(1+100t)g(t)-\epsilon(1+100t)t{\rm R}(g(t))g(t)$ for all $t\in[0,T]\cap[0,\frac{1}{200}]$
@@
    $\sec(g(0))\geq-\epsilon$
    $\implies$ $\sec(g(t))\geq-\epsilon(1+100t)-\epsilon(1+100t)t{\rm R}(g(t))$ for all $t\in[0,T]\cap[0,\frac{1}{200}]$
@@
The proof of the Lemma involves examining the evolution of Tensors/ geometric inequalities under the Ricci flow and using the Maximum principle of [Ham82, Ham86]. For example in case (c), on a closed manifold, one first writes the evolution equation of the tensor $T(\cdot,t)={\rm Ric}(g(t))+\epsilon(1+100t)g(t)+\epsilon(1+100t)t{\rm R}(g(t))g(t)$ as $\frac{\partial}{\partial t}T=\Delta_{g}T+L$ : with the help of the evolution equations for the Ricci curvature and scalar curvature computed in [Ham82], one can explicitly calculate $L$. Then, using the Maximum Principle of Hamilton, one shows that $T(\cdot,0)>0\implies T(\cdot,t)>0.$ The term in $L$ coming from $(\frac{\partial}{\partial t}-\Delta_{g(t)})(t{\rm R}(g(t))),$ is a non-negative term, and plays an important role in the maximum principle argument : It guarantees that $L$ will be positive at any potential first time $t_{0}>0$ and point $p_{0}$ and non-zero direction $v_{0}$ where $T(p_{0},t_{0})(v_{0},v_{0})=0$$\Box$.□ .
@@
Assume for some $\epsilon\in[0,\frac{1}{100}]$ that ${\rm Ric}(g(0))\geq-\epsilon g(0)$ $(\sec(g(0))\geq-\epsilon)$ and that the solution satisfies $|{\rm R}(g(t))|\leq\frac{c^{2}_{0}}{t}.$ Then ${\rm Ric}(g(t))\geq-\epsilon(1+c^{2}_{0})(1+100t)g(t)$ $(\sec(g(t))\geq-\epsilon(1+c^{2}_{0})(1+100t))$
@@
iv) In the case that we have merely a negative-bound from below at time zero, for example $\mathcal{R}(g(0))+kI\geq 0$ for a $k>0$, attempts to generalise the above ”Ingredient 2” to higher dimensions by adding multiples of $t{\rm R}(g(t))I$ and $I$ to $\mathcal{R}(g(t))$ for example, have, to this date, not been successful, except in the case that one already knows $|{\rm Rm}(g(t))|\leq\frac{c^{2}_{0}}{t}$ and $c^{2}_{0}\leq\epsilon(n),$ with $\epsilon(n)$ sufficiently small, as explained in Section 3. A break through for the case that the bound from below is negative in higher dimensions was achieved by Bamler/Cabezas/Rivas [BCRW19]. There they examine the quantity $\ell(t)=\inf\{\alpha\in[0,\infty)\ |\ \mathcal{R}(g(t))+\alpha I\in\mathcal{C}\}$ where $\mathcal{C}$ represents the curvature condition (cone) being considered. Due to some scaling arguments (as in the proof of Theorem 4.5 above), it suffices to consider the case that one has at time zero $\ell(0)\leq\epsilon(n,v_{0}),$ where $\epsilon(n,v_{0})>0$ is as small as we like. They show, that the reaction equation of $\ell(t)$ satisfies an inequality of the type
@@
in the setting that $|{\rm Rm}(g(t))|\leq\frac{c^{2}_{0}}{t}$ which then allows them to obtain a bound on $\ell(t)$ by comparing $\ell$ to the solution $u$ to (4.3) whose initial value is constant, $u(0):=\epsilon$.
@@
In some cases the initial manifold being considered may have some bound on curvature from below, but no bound on the norm of the curvature. Hence even starting the Ricci flow presents difficulties. We survey here some results on the existence of a Ricci flow in this setting. As a prototype case, we consider a three dimensional smooth, complete Riemannian manifold $(M^{3},g(0))$ which is uniformly non-collapsed, that is $V(B_{g(t)}(x,1))\geq v_{0}>0$ for all $x\in M,$ ${\rm Ric}(g(0))\geq-\epsilon,$ $\epsilon\leq\epsilon(v_{0}),$ but we do not assume bounded curvature. That is $\sup_{M}|{\rm Rm}(g(0))|=\infty$ is possible.
@@
In the process of producing the local Ricci flow, it was noticed that the arguments are purely local. These methods were subsequently pushed even further to obtain so called Pyramid Ricci flows, $g(t)$ which are defined on space time regions $\Omega\subseteq M\times(0,T),$ $\Omega:=\cup_{i=1}^{\infty}B_{g(0)}(p,i)\times[0,T_{i}]$ where $T_{1}(v_{0},k)>T_{2}(v_{0},k)>T_{3}(v_{0},k)\ldots$ and $T_{i}\to 0$ as $i\to\infty,$ under the assumption that $V_{g(0)}(B_{g(0)}(p,1))\geq v_{0},$ and $\mathcal{R}(g(0))+kI\in\mathcal{C}$ where $\mathcal{C}$ is one of the conditions (i)-(iv) or (vi)-(ix), and $p\in M$ is a fixed point. As an application, one can prove for example, that any possible Gromov-Hausdorff pointed limit of smooth three manifolds $(M^{3}_{i},g_{i}(0),p_{i})$ satisfying $V_{g_{i}(0)}(B_{g_{i}(0)}(p_{i},1))\geq v_{0},$ and ${\rm Ric}(g_{i}(0))+kg_{i}(0)\geq 0$ is a topological manifold : Simon/Topping [ST21]. See the following papers for further results of this type and further information on this setting: [ST21], McLeod/Topping [MT22], McLeod/Topping [MT20], McLeod [Mcl18], [Hoc19], [LT21b], (cf. Liu [Liu18],[Liu15]).
@@
(Chan/Chen/Lee, Theorem 1.4 in [CCL22]) For any positive integer $n\geq 3$ and constant $A\geq 1000n$, there exists a constant $\epsilon_{0}(n,A)>0$ such that the following holds. Suppose $(M_{i}^{n},g_{i}(0),p_{i})$ is a pointed sequence of smooth, complete, connected Riemannian manifolds with bounded curvature satisfying the following properties:
@@
where $\nu$ is the localised entropy, a localisation of Perleman’s entropy introduced in [Wan18]. Then there exists solutions $(M_{i},g_{i}(t)),p_{i})_{t\in[0,T)}$ for a uniform $T>0$ (independent of $i$) satisfying, $|{\rm Rm}(g_{i}(t)|\leq\frac{C(n,A)\epsilon_{0}}{t},$ $inj(M_{i},g_{i}(t))\geq\gamma(n)\sqrt{t}$ for all $t\in(0,T)$ and, after taking a subsequence, the pointed Cheeger-Hamilton limit $(M,g(t),p)_{t\in(0,T)}$ of $(M_{i},g_{i}(t)),p_{i})_{t\in(0,T)}$ exists and also satisfies $|{\rm Rm}(g(t)|\leq\frac{C(n,A)\epsilon_{0}}{t},$ $inj(M,g(t))\geq\gamma(n)\sqrt{t}$ and
$(\int_{B_{g(t)}(q,2)}|{\rm Rm}(g(t)|^{\frac{n}{2}}dV_{g(t)})^{\frac{2}{n}}\leq C(n,A)\epsilon_{0},$